# A Bimodal Metallicity Distribution Function in the Ultra-Faint Dwarf Galaxy Reticulum II

### v1

@@ 1 INTRODUCTION
In the CDM model, massive galaxies such as the Milky Way have many small satellite galaxies due to hierarchical structure formation (Springel et al., 2008). The decades-long “missing satellite problem” (Klypin et al., 1999; Moore et al., 1999; Simon & Geha, 2007; Brooks et al., 2013; Kim et al., 2018) in which too few satellites were observed was largely resolved as fainter systems were discovered in the era of digital photometric surveys, such as the Sloan Digital Sky Survey (SDSS, Willman et al. 2005; Zucker et al. 2006; Belokurov et al. 2008, 2009, 2010; Kim et al. 2015) and the Dark Energy Survey (DES, Bechtol et al. 2015; Drlica-Wagner et al. 2015, 2016; Kim & Jerjen 2015; Koposov et al. 2015a, 2018; Martin et al. 2015; Torrealba et al. 2016). The faintest newly discovered satellites became known as ultra-faint dwarf galaxies (UFDs), which are galaxies that live in the smallest dark matter halos and are extremely dark matter dominated (Simon & Geha, 2007; Willman & Strader, 2012; Simon, 2019). They are the most pristine systems, with little to no chemical evolution, due to the majority of their star formation occurring before it is thought that they were quenched during the epoch of reionization (Bullock et al., 2000; Benson et al., 2002; Somerville, 2002; Brown et al., 2014; Rodriguez Wimberly et al., 2019). They may also have been important during reionization by contributing ionizing UV photons (Weisz & Boylan-Kolchin, 2017; Wu & Kravtsov, 2024).
@@
UFDs contain some of the oldest and most metal-poor () stars since they formed their stars during the early Universe when there were little to no metals in the gas available for star formation (Simon, 2019). Therefore, UFDs are relics of the first galaxies formed in the early Universe (Bovill & Ricotti, 2009, 2011). This makes them great sites to search for the signatures of the first stars in the Universe, such as enhanced carbon abundances (Frebel & Bromm, 2012; Frebel & Norris, 2015; Jeon et al., 2017).
@@
The metallicity distribution function (MDF) is the distribution of individual stellar metallicities within a galaxy. The distribution is determined by a galaxy’s star formation history and is affected by stellar yields, feedback, and star formation efficiency. MDFs have been extensively studied in classical dwarf spheroidals (dSphs, ) such as in Carigi et al. (2002); Lanfranchi et al. (2008); Tolstoy et al. (2009); Kirby et al. (2009, 2011); Norris et al. (2010). The MDFs are fit with analytic chemical evolution models (e.g., Lynden-Bell 1975; Pagel 1997; Lanfranchi & Matteucci 2003; Johnson et al. 2021; Sandford et al. 2024) to infer their gas inflows/outflows. These classical dwarf satellites are sufficiently massive that a well-populated MDF can be determined only using metallicities of red giant branch (RGB) stars.
@@
Pushing down to the ultra-faint regime (), there are limited metallicity measurements as these galaxies contain fewer RGB stars. Simon & Geha (2007) published the first metallicity measurements in UFDs, but obtaining larger UFD MDF samples still remains an ongoing area of research (e.g., Norris et al., 2010; Lai et al., 2011; Vargas et al., 2013; Brown et al., 2014; Romano et al., 2015; Simon et al., 2015; Jenkins et al., 2021; Fu et al., 2022, 2023). The current best-studied UFD MDF is of Boötes I, which had an MDF of 16 stars immediately after its discovery and soon increased to 41 stars (Norris et al., 2010; Lai et al., 2011). A decade later, its most recent MDF consists of 70 stars, the largest spectroscopic MDF of any UFD (Jenkins et al., 2021). Another way to get more populated MDFs is using photometric metallicities, which can be fairly precise by using narrowband filters (Longeard et al., 2018; Chiti et al., 2020; Fu et al., 2022, 2023). With deep narrowband CaHK HST photometry, Fu et al. (2022) obtained metallicities of 60 RGB stars in Eridanus II, noting that their results are consistent with spectroscopic studies (Li et al., 2017), but with larger scatter.
@@
In this paper, we study the MDF of Reticulum II (Ret II), an ultra-faint dwarf galaxy with stellar mass of (Mutlu-Pakdil et al., 2018). Ret II was discovered from the first data release of the Dark Energy Survey (DES) in 2015 (Bechtol et al., 2015; Koposov et al., 2015a) and quickly confirmed to be a galaxy (Simon et al., 2015; Walker et al., 2015; Koposov et al., 2015b). As one of the closest UFDs discovered at only 32 kpc, it became a galaxy of interest due to the feasibility of deeper studies. It was initially speculated to have a connection to the Magellanic Clouds (MCs), which was later confirmed with orbital histories (Patel et al., 2020; Erkal & Belokurov, 2020). It was also of interest because of the tentative detection of gamma-ray emission (e.g., Geringer-Sameth et al. 2015), which could be due to dark matter particles self-annihilating into high-energy gamma rays. With high-resolution follow-up spectroscopy, Ret II was discovered to be the first UFD to have been enriched by a rare and prolific r-process nucleosynthesis event, either a neutron star merger or a rare core-collapse supernova (Ji et al., 2016a; Roederer et al., 2016). More recently, several works have studied the detailed formation history of Ret II. The star formation history (SFH) described in Simon et al. (2023) favors a two-burst star formation scenario. Ji et al. (2023) details the low dispersion in the r-process enriched stars. Lastly, Fu et al. (2023) obtained a photometric MDF from 76 metallicities (50 constrained and 26 upper limits).
| Mask | Observing Date | R.A. | Decl. | Num. of slits | Useful spectra* | Num. of members | |
| (h:m:s) | (d:m:s) | (s) | (%) | ||||
| retIIa | 25 Nov 2019 | 03:35:30 | 54:02:50 | 21600 | 126 | 89% | 94 |
| retIIb | 26 Nov 2019 | 03:35:25 | 54:03:00 | 24800 | 122 | 84% | 82 |
| retIIc | 17 Nov 2023 | 03:35:37 | 54:01:50 | 20640 | 99 | 71% | 62 |
Note. — * After removing galaxy spectra and SNR 3.
@@
Ret II has only 20 stars with spectroscopic metallicities in the literature, excluding upper limits (Simon et al., 2015; Ji et al., 2016a, 2023). This paper aims to determine a well-populated spectroscopic MDF. Due to the small number of RGB stars, we had to determine metallicities of the faint main sequence turn-off (MSTO) stars (g 22), the first time this has been done in any dwarf spheroidal fainter than the Magellanic Clouds.
@@ 2 OBSERVATIONS
We observed stars in Reticulum II with the Inamori-Magellan Areal Camera and Spectrograph (IMACS; Dressler et al. 2011) with the f/2 camera on the Magellan Baade telescope at Las Campanas Observatory on 2019 November 26 and 27 with follow-up observations on 2023 November 17. The targets were selected in 3 ways: (1) along the isochrones of the deep () photometric catalogs provided by Mutlu-Pakdil et al. (2018) using Megacam (McLeod et al., 2015) on the Magellan Clay Telescope, (2) similarly from Simon et al. (2023) using Advanced Camera for Surveys (ACS; Ford et al. 2003) on the Hubble Space Telescope, and (3) from a sample of proper motion members in Pace & Li (2019) using DES DR1 (Bechtol et al., 2015; Drlica-Wagner et al., 2018; Abbott et al., 2018) crossmatched with Gaia DR2 (Gaia Collaboration et al., 2018). We obtained low-resolution spectroscopy (R1500 near the blue end of the spectrum) using the 400 /mm grism with a central wavelength of Å. The 3 observing nights described in Table 1 each used different multi-slit masks with 1” slits, totaling 245 unique objects. The wavelength range for each spectrum spans Å. Each slitmask had 6 hours of total exposure time allowing us to reach signal-to-noise (SNR) 10 per 0.9Å pixel near CaK (3933Å) for stars near the MSTO of Reticulum II at g 21. Figure 1 shows the fraction of photometric candidates observed per magnitude bin. For the MSTO stars, we have completeness.
We followed Newman et al. (2020) using the CarPy (Kelson et al., 2000; Kelson, 2003) pipeline for data reduction. The following procedure is run per slitmask. A mask image is taken during the afternoon that maps the positions of the slits to observed objects. Each science frame has a corresponding HeHg arc lamp spectrum for wavelength calibration and finding the object position. The pipeline uses the [OI] 5577Å sky emission line to tweak the wavelength solution before sky subtraction. Flat-field twilight spectra are used to divide out pixel-to-pixel variations of the eight CCD detectors. CarPy provides the 2D wavelength calibrated rectified spectra. The one-dimensional spectrum is then extracted using the optimal method described in Horne (1986) that traces the stellar profile to apply larger weights to pixels that contribute more flux. On the final observing night, issues with the Atmospheric Dispersion Corrector (ADC) resulted in distortions in the spatial direction. We thus added an additional correction by fitting out this spatial distortion.
@@ 3.1 Membership selection
We observed spectra of 245 objects from photometric catalogs of Ret II in the literature. After cleaning out 7 clear galaxy/quasar contaminants and 7 other non-stellar spectra, we made further membership selections to reject possible Milky Way contaminants. We required the data to have SNR 3 to ensure good data quality. Some stars were observed over multiple nights. For these stars, the total flux of a star is the sum of the fluxes of each night, and the total uncertainty is the quadrature sum of the uncertainties. We selected members along a 13 Gyr isochrone with using MESA Isochrones and Stellar Tracks (MIST; Dotter 2016; Choi et al. 2016; Paxton et al. 2011, 2013, 2015) and Minimint (MIni Mist INTerpolation; Koposov 2023). We used a distance modulus of 17.5 and reddening of 0.04, within the uncertainty of 0.01 mag measured in Simon et al. (2023). We rejected 28 outliers with a distance from the isochrone of 0.08. We had a subset of 50 stars with proper motions from Gaia DR3, and rejected 8 stars larger than 3 from the mean proper motion of Ret II found in Pace et al. (2022). Figure 2 shows the membership selection criteria. We also rejected 7 stars that Simon et al. (2015) labeled as nonmembers based on their velocities.
We found radial velocities by cross-correlating the stellar spectra with the high-resolution rest-frame template spectrum of HD21581 around the H (4340Å) and H (4861Å) lines. HD21581 is an RGB star with a heliocentric radial velocity of 153.7 km s-1 and , a higher metallicity than the average metallicity in Reticulum II (Roederer et al., 2014). The overlapping wavelength range between our spectra and HD21581 is 3800 to 5550Å. We used the strong absorption lines near the redder end of the spectrum to avoid the lines at the noisier blue end and minimize atmospheric dispersion effects. Only one star (HST_037) did not have a velocity because the H and H lines were near the chip gaps in the detector, but it passed the velocity cut with the Ca K line velocity. We show the velocity histogram in Figure 2. We rejected 21 stars with velocities larger than 125 km s-1, too large to be consistent with the mean velocity of Reticulum II. Given the low resolution, the velocity uncertainties are km s-1, which is not useful for the internal kinematics of the galaxy. Implementing all membership selection criteria resulted in a sample of 167 members.
@@ 3.2 Measuring [Fe/H]
We measured metallicities from a calibration (Beers et al., 1999) using a pseudo-equivalent width (EW) of the Ca II K line (3933Å) and color index as input parameters. The EW measurement followed the KP Index method described in the calibration paper and outlined in Chiti et al. (2018). We calculated the area under the Ca II K line within 6Å, 12Å, and 18Å bands centered at 3933.7Å. Figure 3 shows the spectral range used for the analysis (K6 from Å, K12 from Å, and K18 from Å). The resulting equivalent width adopts the KP index value based on the following criteria:
@@
where is the dispersion of Å per pixel, is the normalized flux uncertainity at wavelength, , and is the length of the KP region. We report the EW and values in Appendix A. We fit a Gaussian profile to the Ca K region for each star to check for line detections. The line fit for each star was inspected by eye to confirm whether there was a real detection. We determined 129 detections and 38 non-detections. We show the EW and values colored by metallicity in Figure 4. The typical EW uncertainties are 0.29, 0.43, 0.19, 0.15, and 0.09Å for the following color bins: . For non-detections, we calculated upper limits. The are propagated into the metallicity uncertainties, . The majority of stars were dominated by the systematic uncertainty in the calibration, . In Beers et al. (1999), there were not many stars in their sample near the low KP and low regions, leading to uncertainties of .
We used Magellan/Megacam photometry from Mutlu-Pakdil et al. (2018) for the input. After dereddening, and -band magnitudes were transformed to SDSS magnitudes (Drlica-Wagner et al., 2018) then to UBVRI (Jordi et al., 2006). Figure 5 shows the color-magnitude diagram of the member stars colored by their metallicities. We observed 10 stars without Magellan/Megacam photometry. From 150 stars with Magellan/Megacam, DES, and HST magnitudes, we compared the magnitude systems and determined the coefficients for each magnitude relation. We used this to transform DES and HST magnitudes onto a scale similar to the Magellan/Megacam magnitudes. The scatter of the relation ( mag) was propagated in the photometric uncertainty, , for the 10 stars. The total metallicity uncertainty for a star is .

### v2

@@ 1 INTRODUCTION
In the CDM model, massive galaxies such as the Milky Way have many small satellite galaxies due to hierarchical structure formation (Springel2008). The decades-long “missing satellite problem” (Klypin1999; Moore1999; Simon&Geha2007; Brooks2013; Kim2018) in which too few satellites were observed was largely resolved as fainter systems were discovered in the era of digital photometric surveys, such as the Sloan Digital Sky Survey (SDSS, Willman2005; Zucker2006; Belokurov2008; Belokurov2009; Belokurov2010; Kim2015) and the Dark Energy Survey (DES, Bechtol+2015; Drlica-Wagner2015; Drlica-Wagner2016; Kim&Jerjen2015; Koposov2015a; Koposov2018; Martin2015; Torrealba2016). The faintest newly discovered satellites became known as ultra-faint dwarf galaxies (UFDs), which are galaxies that live in the smallest dark matter halos and are extremely dark matter dominated (Simon&Geha2007; Willman2012; Simon2019). They are the most pristine systems, with little to no chemical evolution, due to the majority of their star formation occurring before it is thought that they were quenched during the epoch of reionization (Bullock2000; Benson2002; Somerville2002; Brown2014; RodriguezWimberly2019). They may also have been important during reionization by contributing ionizing UV photons (Weisz2017; WuKravtsov2024).
@@
UFDs contain some of the oldest and most metal-poor () stars since they formed their stars during the early Universe when there were little to no metals in the gas available for star formation (Simon2019). Therefore, UFDs are relics of the first galaxies formed in the early Universe (Bovill2009; Bovill2011). This makes them great sites to search for the signatures of the first stars in the Universe, such as enhanced carbon abundances (Frebel&Bromm2012; Frebel&Norris2015; Jeon2017).
@@
The metallicity distribution function (MDF) is the distribution of individual stellar metallicities within a galaxy. The distribution is determined by a galaxy’s star formation history and is affected by stellar yields, feedback, and star formation efficiency. MDFs have been extensively studied in classical dwarf spheroidals (dSphs, ) such as in Carigi2002; Lanfranchi2008; Tolstoy2009; Kirby2009; Kirby+2011; Norris2010. The MDFs are fit with analytic chemical evolution models (e.g., Lynden-Bell1975; Pagel1997; Lanfranchi2003; Johnson2021; Sandford2024) to infer their gas inflows/outflows. These classical dwarf satellites are sufficiently massive that a well-populated MDF can be determined only using metallicities of red giant branch (RGB) stars.
cccccrrr
[…]
(h:m:s) (d:m:s) (s) (%)
[…]
retIIb 26 Nov 2019 03:35:25 54:03:00 24800 122 84% 82
retIIc 17 Nov 2023 03:35:37 54:01:50 20640 99 71% 62
[…]
@@
Pushing down to the ultra-faint regime (), there are limited metallicity measurements as these galaxies contain fewer RGB stars. Simon&Geha2007 published the first metallicity measurements in UFDs, but obtaining larger UFD MDF samples still remains an ongoing area of research (e.g., Norris2010; Lai2011; Vargas2013; Brown2014; Romano2015; Simon+2015; Jenkins2021; Fu+2022; Fu+2023). The current best-studied UFD MDF is of Boötes I, which had an MDF of 16 stars immediately after its discovery and soon increased to 41 stars (Norris2010; Lai2011). A decade later, its most recent MDF consists of 70 stars, the largest spectroscopic MDF of any UFD (Jenkins2021). Another way to get more populated MDFs is using photometric metallicities, which can be fairly precise by using narrowband filters (Longeard2018; Chiti2020; Fu+2022; Fu+2023). With deep narrowband CaHK HST photometry, Fu+2022 obtained metallicities of 60 RGB stars in Eridanus II, noting that their results are consistent with spectroscopic studies (Li2017), but with a larger scatter.
@@
In this paper, we study the MDF of Reticulum II (Ret II), an ultra-faint dwarf galaxy with stellar mass of (Mutlu-Pakdil+18). Ret II was discovered from the first data release of the Dark Energy Survey (DES) in 2015 (Bechtol+2015; Koposov2015a) and quickly confirmed to be a galaxy (Simon+2015; Walker2015; Koposov2015b). As one of the closest UFDs discovered at only 32 kpc, it became a galaxy of interest due to the feasibility of deeper studies. It was initially speculated to have a connection to the Magellanic Clouds (MCs), which was later confirmed with orbital histories (Patel2020; Erkal2020). It was also of interest because of the tentative detection of gamma-ray emission (e.g., GeringerSameth2015), which could be due to dark matter particles self-annihilating into high-energy gamma rays. With high-resolution follow-up spectroscopy, Ret II was the first UFD discovered to have been enriched by a rare and prolific r-process nucleosynthesis event, either a neutron star merger or a rare core-collapse supernova (Ji+2016a; Roederer2016). More recently, several works have studied the detailed formation history of Ret II. The star formation history (SFH) described in Simon+2023 favors a two-burst star formation scenario. Ji+2023 details the low dispersion in the r-process enriched stars. Lastly, Fu+2023 obtained a photometric MDF from 76 metallicities (50 constrained and 26 upper limits).
@@
Ret II has only 20 stars with spectroscopic metallicities in the literature, excluding upper limits (Simon+2015; Ji+2016a; Ji+2023). This paper aims to determine a well-populated spectroscopic MDF. Due to the small number of RGB stars, we had to determine metallicities of the faint main sequence turn-off (MSTO) stars (g 22), the first time this has been done in any dwarf spheroidal fainter than the Magellanic Clouds.
@@ 2 OBSERVATIONS
We observed stars in Reticulum II with the Inamori-Magellan Areal Camera and Spectrograph (IMACS; Dressler+2011) with the f/2 camera on the Magellan Baade telescope at Las Campanas Observatory on 2019 November 26 and 27 with follow-up observations on 2023 November 17. The targets were selected in 3 ways: (1) along the isochrones of the deep () photometric catalogs provided by Mutlu-Pakdil+18 using Megacam (McLeod2015) on the Magellan Clay Telescope, (2) similarly from Simon+2023 using Advanced Camera for Surveys (ACS; Ford+2003) on the Hubble Space Telescope, and (3) from a sample of proper motion members in Pace&Li+19 using DES DR1 (Bechtol+2015; Drlica-Wagner2018; Abbott2018) crossmatched with Gaia DR2 (GaiaCollab+2018a:Brown). We obtained low-resolution spectroscopy (R1500 near the blue end of the spectrum) using the 400 /mm grism with a central wavelength of Å. The 3 observing nights described in Table 1 each used different multi-slit masks with 1” slits, totaling 245 unique objects. The wavelength range for each spectrum spans Å. Each slitmask had 6 hours of total exposure time allowing us to reach signal-to-noise (SNR) 10 per 0.9Å pixel near CaK (3933Å) for stars near the MSTO of Reticulum II at g 21. In Figure 1, the left panel shows the CMD of stars in the Mutlu-Pakdil+18 catalog in gray circles and the stars selected for spectroscopic follow-up in green circles, and the right panel shows the fraction of photometric candidates from Mutlu-Pakdil+18 observed per magnitude bin. For the MSTO stars, we have completeness.
We followed Newman+2020 using the CarPy (Kelson+2000; Kelson2003) pipeline for data reduction. The following procedure is run per slitmask. A mask image is taken during the afternoon that maps the positions of the slits to observed objects. Each science frame has a corresponding HeHg arc lamp spectrum for wavelength calibration and finding the object position. The pipeline uses the [OI] 5577Å sky emission line to tweak the wavelength solution before sky subtraction. Flat-field twilight spectra are used to divide out pixel-to-pixel variations of the eight CCD detectors. CarPy provides the 2D wavelength calibrated rectified spectra. The one-dimensional spectrum is then extracted using the optimal method described in Horne1986 that traces the stellar profile to apply larger weights to pixels that contribute more flux. On the final observing night, issues with the Atmospheric Dispersion Corrector (ADC) resulted in distortions in the spatial direction. We thus added an additional correction by fitting out this spatial distortion.
@@ 3.1 Membership selection
We observed spectra of 245 objects from photometric catalogs of Ret II in the literature. After cleaning out 7 clear galaxy/quasar contaminants and 7 other non-stellar spectra, we made further membership selections to reject possible Milky Way contaminants. We required the data to have SNR 3 to ensure good data quality. Some stars were observed over multiple nights. For these stars, the total flux of a star is the sum of the fluxes of each night, and the total uncertainty is the quadrature sum of the uncertainties. We selected members along a 13 Gyr isochrone with using MESA Isochrones and Stellar Tracks (MIST; Dotter2016; Choi2016; Paxton2011; Paxton2013; Paxton2015) and Minimint (MIni Mist INTerpolation; sergey_koposov_2023_10437032). We used a distance modulus of 17.5 and reddening of 0.04, within the uncertainty of 0.01 mag measured in Simon+2023. We rejected 28 outliers with a distance from the isochrone of 0.08. We had a subset of 50 stars with proper motions from Gaia DR3, and rejected 8 stars larger than 3 from the mean proper motion of Ret II found in Pace2022. Figure 2 shows the membership selection criteria. We also rejected 7 stars that Simon+2015 labeled as nonmembers based on their velocities.
We found radial velocities by cross-correlating the stellar spectra with the high-resolution rest-frame template spectrum of HD21581 around the H (4340Å) and H (4861Å) lines. HD21581 is an RGB star with a heliocentric radial velocity of 153.7 km s-1 and , a higher metallicity than the average metallicity in Reticulum II (Roederer2014). The overlapping wavelength range between our spectra and HD21581 is 3800 to 5550Å. We used the strong absorption lines near the redder end of the spectrum to avoid the lines at the noisier blue end and minimize atmospheric dispersion effects. Only one star (HST_037) did not have a velocity because the H and H lines were near the chip gaps in the detector, but it passed the velocity cut with the Ca K line velocity. We show the velocity histogram in Figure 2. We rejected 21 stars with velocities larger than 125 km s-1, too large to be consistent with the mean velocity of Reticulum II. Given the low resolution, the velocity uncertainties are km s-1, which is not useful for the internal kinematics of the galaxy. Implementing all membership selection criteria resulted in a sample of 167 members.
@@ 3.2 Measuring [Fe/H]
We measured metallicities from a calibration (Beers+1999) using a pseudo-equivalent width (EW) of the Ca II K line (3933Å) and color index as input parameters. The EW measurement followed the KP Index method described in the calibration paper and outlined in Chiti2018. We calculated the area under the Ca II K line within 6Å, 12Å, and 18Å bands centered at 3933.7Å. Figure 3 shows the spectral range used for the analysis (K6 from Å, K12 from Å, and K18 from Å). The resulting equivalent width adopts the KP index value based on the following criteria:
@@
where is the dispersion of the 400 l/mm grism Å per pixel, is the normalized flux uncertainity at wavelength, , and is the length of the KP region. We report the EW and values in Appendix A. We fit a Gaussian profile to the Ca II K line for each star to check for line detections. The line fit for each star was inspected by eye to confirm whether there was a real detection. We determined 129 detections and 38 non-detections. We show the EW and values colored by metallicity in Figure 4. The typical EW uncertainties are 0.29, 0.43, 0.19, 0.15, and 0.09Å for the following color bins: . For non-detections, we calculated upper limits. The are propagated into the metallicity uncertainties, . The majority of stars were dominated by the systematic uncertainty in the calibration, . In Beers+1999, there were not many stars in their sample near the low KP and low regions, leading to uncertainties of . The color does not appear to have an effect on the systematic uncertainties, while higher equivalent widths have lower uncertainties.
We used Magellan/Megacam photometry from Mutlu-Pakdil+18 for the input. After dereddening, and -band magnitudes were transformed to SDSS magnitudes (Drlica-Wagner2018) then to UBVRI (Jordi+2006). Figure 5 shows the color-magnitude diagram of the member stars colored by their metallicities. We observed 10 stars without Magellan/Megacam photometry. From 150 stars with Magellan/Megacam, DES, and HST magnitudes, we compared the magnitude systems and determined the coefficients for each magnitude relation. We used this to transform DES and HST magnitudes onto a scale similar to the Magellan/Megacam magnitudes. The scatter of the relation ( mag) was propagated in the photometric uncertainty, , for the 10 stars. The total metallicity uncertainty for a star is .